# Estimating the Probability of Sampling a Trained Neural Network at Random

### v1

@@ 3.3 Gaussian volume
Note that the integrand is of the form $\exp(\text{quadratic}(r)+n\log r)$ and varies rapidly when $n$ is large. We evaluate these integrals numerically using an approximation similar to Lagrange’s method, expanding the exponent to second-order and performing a Gaussian integral using a numerically-stable implementation of the error function. In practice, the error from the approximation is less than floating-point rounding error.
3.4 From loss to KL divergence
The forms of the volume hypothesis above deal with the training loss. In practice, however, the Hessian of the loss function is closely related to the Fisher information matrix for the model (Martens & Grosse, 2015), which is the Hessian of KL divergence. We therefore use KL from the anchor point as our cost function, as well as our defining criterion for our neighborhoods. This has the added benefit of putting the anchor point at a global minimum by definition, simplifying some of the practical challenges. It can also be interpreted within an MDL framework as the description length of the network, as shown below.
@@ 4.1 Preconditioners
We also use the second-moment Adam preconditioner for Pythia and ConvNeXt, where it shows both a clear improvement in the value of the estimates and a smaller sample variance (Figures 2 and 3). The improvement is several standard deviations above most of the naïve estimates, suggesting that it would be infeasible to merely increase the sample size to try to get the same result. [^7] Note, however, that the naïve estimate for the unpoisoned ConvNeXt network has a large outlier sample that completely dominates the aggregated estimate, nearly reaching the bulk of the preconditioned estimates.
For ConvNeXt, we find that the poisoned network has a smaller local volume (with or without preconditioning), in agreement with the results of Huang et al. (2020) on small networks and in line with our expectation from the MDL and compression perspective.
@@ 4.3 Across cutoffs
When varying the cutoff, we see a roughly power-law trend for local volume on Pythia (6). The log-log slope is consistent with $n/2$, where $n$ is the model dimension; this is what would be expected for a purely quadratic cost function. A line with this slope is shown in black for comparison. At very high cutoffs ($\geq 10$ nats), the Adam preconditioner begins to fail, although raising ${\epsilon}$ partially counteracts this.
For ConvNeXt, the result is similar for cutoffs between $10^{-5}$ and $10^{-2}$. The preconditioner again fails at high cutoff, sooner than for Pythia. At very low cutoffs ($10^{-6}$ nats), the poisoned network’s local volume suddenly plummets. We have not investigated this phenomenon in detail, but we suspect it may have to do with floating point rounding error.
The slope of $n/2$ is somewhat in conflict with empirical results for the Local Learning Coefficient (Hoogland et al., 2024; Wang et al., 2024), which we suspect may be an artifact of the Monte Carlo local volume estimator. In particular, at low $k$, our estimator is nearly insensitive to very flat (long) directions, unless they are adequately corrected by the preconditioner. If the steeper (shorter) directions are close to quadratic in shape, this could cause the cost function to appear more quadratic than it really is.
@@ 5 Conclusion
Our results are broadly consistent with a weak form of the volume hypothesis. As expected, poisoned networks were observed to have smaller local volumes than unpoisoned ones. That said, more research is needed to confirm or refute any specific version of the volume hypothesis.
One promising direction for future work may be to use stochastic gradient Langevin dynamics (SGLD) to propose directions along which to measure the neighborhood’s radius. We also are excited to see practical applications of local volume estimation. We think it may be useful for predicting generalization performance. More speculatively, if we define the cost function to be the model’s behavior on a relatively narrow distribution– say, a set of math problems fed to a large language model– the local volume may tell us something about how “difficult” these problems are for the model, or how hard it is “thinking.”
Another possible direction may be to estimate the probability measure of neighborhoods around initializations that lead to a given final behavior after training, which corresponds almost exactly to the probability of SGD producing that trained behavior. This would allow for a precise quantitative evaluation of the basin volume hypothesis, and could potentially be accomplished via the training Jacobian (Belrose & Scherlis, 2024).
@@
Adam Scherlis came up with the main ideas of this paper, wrote the code and performed all experiments and data analysis, and did much of the writing. Nora Belrose proposed the idea of using expected KL divergence as a cost function in lieu of training loss, pointed out the connection with minimum description length, proposed a last-minute reframing of the paper in terms of “sampling trained networks at random,” and wrote various parts of the paper.
@@
Appendix A Variance of the log-estimator
@@
If the spectrum of $\mathbf{A}$ has any significant variance, as is empirically often the case for neural network Hessians, this variance will be extremely large.

### v2

@@ 5.3 Gaussian volume
Note that the integrand is of the form $\exp(\text{quadratic}(r)+n\log r)$ and varies rapidly when $n$ is large.
We evaluate these integrals numerically using an approximation similar to Lagrange’s method, expanding the exponent to second-order and performing a high-dimensional Gaussian integral using a numerically-stable implementation of the error function. In practice, the error from the approximation is less than floating-point rounding error.
5.4 Choice of KL over loss for neighborhoods
The forms of the volume hypothesis above deal with the training loss, and some, such as Appendix B, have an obvious relationship to training-loss neighborhoods. On the other hand, KL neighborhoods are more compact, more naturally captured by the notion of a star domain, and have a cost minimum at the anchor point. In addition, the Hessian of the loss function is closely related to the Fisher information matrix for the model (Martens & Grosse, 2015), which is the Hessian of KL divergence; this leads us to believe that experiments on KL neighborhoods are a good proxy for loss neighborhoods, as far as testing the volume hypothesis goes. We therefore choose KL local volume as the target for our experiments.
@@ 6 Results
For ConvNeXt on CIFAR-10, we also report results on the clean split of the training data in Appendix D. This is the best test of our estimator as a practical interpretability tool: it shows that we can detect poor generalization (equivalently, excess complexity) in the poisoned model using a small number of forward passes over training data on which the model’s behavior is indistinguishable from the well-generalizing unpoisoned model.
@@ 6.1 Preconditioners
We also use the second-moment Adam preconditioner for Pythia and ConvNeXt, where it shows both a clear improvement in the value of the estimates and a smaller sample variance (Figures 2 and 8). The improvement is several standard deviations above most of the naïve estimates, suggesting that it would be infeasible to merely increase the sample size to try to get the same result. [^11] Note, however, that the naïve estimate for the unpoisoned ConvNeXt network has a large outlier sample that completely dominates the aggregated estimate, nearly reaching the bulk of the preconditioned estimates.
For ConvNeXt, we find that the poisoned network has a smaller local volume (with or without preconditioning), in agreement with the results of Huang et al. (2020) on small networks and in line with our expectation from the MDL and compression perspective. Notably, we evaluate local volume on a held-out test set, not the poisoned dataset, demonstrating that the higher network complexity induced by poisoning is visible to our methods even when the poisoned data is unknown.
@@ 6.3 Across cutoffs
When varying the cutoff, we see a roughly power-law trend for local volume on Pythia (6). The log-log slope is consistent with $n/2$, where $n$ is the model dimension; this is what would be expected for a purely quadratic cost function. A line with this slope is shown in black for comparison. At very high cutoffs ($\geq 10$ nats), the Adam preconditioner begins to fail, producing smaller-than-naive estimates, although raising our ${\epsilon}$ hyperparameter counteracts this. (Note that 10 nats corresponds to a perplexity of over 20,000, so this is a fairly extreme cutoff.)
For ConvNeXt, the result is similar for cutoffs between $10^{-5}$ and $10^{-2}$. The preconditioner again fails at high cutoff, sooner than for Pythia. At very low cutoffs ($10^{-6}$ nats), the poisoned network’s local volume suddenly plummets. We have confirmed that this is a floating point precision issue: the radius-finding binary search fails to converge for cutoffs this low.
The consistent slope of $n/2$ for small cutoffs is in contrast with known results on the Local Learning Coefficient (Hoogland et al., 2024; Wang et al., 2024), where this is a strict upper bound that is not seen in practice. We believe that this is due to our choice of KL over the more-singular training loss as a cost function, or our imposition of a Gaussian prior, which (by design) smoothly cuts off the measure of flat directions. [^12] This could be tested by replacing the explicit Gaussian-integral computation with an additional quadratic term in the cost function.
@@ 7 Conclusion
Our results are broadly consistent with the volume hypothesis. As expected, badly-generalizing poisoned networks were observed to have smaller local volumes than well-generalizing unpoisoned ones. That said, more research is needed to confirm or refute any specific version of the volume hypothesis.
One promising direction for future work may be to use stochastic gradient Langevin dynamics (SGLD) to improve the importance-sampling step, by proposing directions along which to measure the neighborhood’s radius.
Another possible application of SGLD is to replace our radius-finding estimator entirely, as is done in Singular Learning Theory research with the LLC estimator. SGLD does not directly give volume information, but thermodynamic integration can be used to compute a volume from a sequence of SGLD equilibria. This would potentially give a more accurate (albeit much slower) local-volume estimator, which could be used as a ground truth to evaluate the accuracy of our estimator.
We are excited to see practical applications of local volume estimation. We think it may be useful for predicting generalization performance. More speculatively, if we define the cost function to be the model’s behavior on a relatively narrow distribution– say, a set of math problems fed to a large language model– the local volume may tell us something about how “difficult” these problems are for the model, or how hard it is “thinking.” We hope that the estimators we describe here can eventually be useful for detecting backdoors, scheming, or other unwanted hidden behavior in models.
Another possible direction may be to estimate the probability measure of neighborhoods around initializations that lead to a given final behavior after training, which corresponds almost exactly to the probability of SGD producing that trained behavior. This would allow for a precise quantitative evaluation of the volume hypothesis, and could potentially be accomplished via the training Jacobian (Belrose & Scherlis, 2024).
@@
Adam Scherlis came up with the main ideas of this paper, wrote the code and performed all experiments and data analysis, and did much of the writing. Nora Belrose proposed the idea of using expected KL divergence as a cost function in lieu of training loss, pointed out the connection with minimum description length, proposed a last-minute reframing of the paper in terms of “sampling trained networks at random,” and wrote various parts of the paper. We thank Jesse Hoogland, Dmitry Vaintrob, and others for useful conversations.
@@
Appendix A Ellipsoidal toy model
In this appendix we collect theoretical results and analysis for the simple case of a quadratic cost function taking its minimum at the anchor point, or (slightly more generally) for an ellipsoidal neighborhood. This case is theoretically tractable and gives qualitatively accurate predictions for actual neighborhoods.
A.1 Variance of the log-estimator
@@
If the spectrum of $\mathbf{A}$ has variance of order $\mathbb{E}[\lambda]^{2}$, as is generally the case for neural network Hessians, this variance will be extremely large (of order $n^{2}$ where $n$ is the parameter count).
A.2 Underestimation and the Hessian spectrum
Consider an ellipsoidal basin, centered at the origin. Suppose the inverse principal axis radii $1/R_{i}$ vary smoothly over several orders of magnitude according to some probability distribution $P(1/R)$. Define a diagonal matrix $H$ s.t.
$H_{ii}=1/{R_{i}^{2}}$ (16)
Then define
$f(x)$ $:=\frac{1}{2}\sum_{ij}H_{ij}x_{i}x_{j}$ (17)
so that $f(x)\leq\frac{1}{2}\iff x\in X$.
Note that the distance $R(\hat{v})$ to the boundary for a unit vector $\hat{v}$ is then given by
$f(R(\hat{v})\hat{v})$ $=\frac{1}{2}$ (18)
$R(\hat{v})^{2}f(\hat{v})$ $=\frac{1}{2}$
$R(\hat{v})^{2}$ $=\frac{1}{\sum_{ij}H_{ij}\hat{v}_{i}\hat{v}_{j}}$
$\frac{1}{R(\hat{v})^{2}}$ $={\sum_{i}\frac{\hat{v}_{i}^{2}}{R_{i}^{2}}}$
If the dimension is high enough, such that $1/R_{i}$ stays nearly constant across $k\gg 1$ consecutive axes, then the $k$ axes closest to some value $1/R$ will contribute to the sum according to
$\sum_{i=n}^{n+k-1}\frac{\hat{v}_{i}^{2}}{R_{i}^{2}}$ $\approx\frac{1}{R^{2}}\sum_{i=n}^{n+k-1}{\hat{v}_{i}^{2}}$ (19)
$\approx\frac{1}{R^{2}}\frac{k}{n}$
where in the second line we have made use of the fact that $N(0,1/n)$ approximates a uniform unit sphere.
Then, summing along the distribution $P(1/R)$, we get
[…]
Therefore, in this limit, $R(\hat{v})^{2}$ will approximate the harmonic mean of $R_{i}^{2}$ with high probability. This gives the wrong volume; the correct volume is $\propto\prod_{i}R_{i}$, so it is a function of the geometric mean of $R_{i}$. (Recall that the harmonic mean is bounded above by the geometric mean.) The naive volume estimator will still be unbiased (as it always is), but most samples will be close to a modal value that is significantly smaller than the true volume. The difference between mean and mode will be produced entirely by a heavy right tail, which would not be evident from a typical sample of a few random vectors.
When applying this analysis to real neural networks, there is an extra complication: some principal axes are outliers, which breaks the law-of-large-numbers approximation used above. However, this actually makes the underestimation problem worse in practice, since the tail (driven by the outliers) will be even heavier.
A.3 Failure of strong volume hypothesis on small scales
If we introduce isotropic noise and solve the resulting Fokker-Planck equation, it can be shown that the log-density instead converges to something proportional to the loss, as in SGLD. However, if the noise is not isotropic – in particular if it is stronger in more steeply-curved directions, as is true in practice – then this fails (Mandt et al., 2018).
Appendix B Another volume hypothesis
Let $\mu_{0}$ be the probability measure on $\mathbb{R}^{N}$ from which the initial network parameters $\theta_{0}$ are sampled, usually a uniform distribution on a compact set or a Gaussian. Let $\mu_{t}$ be the distribution over network parameters at timestep $t$ in training, and let $f_{t}(x)=\frac{d\mu_{t}}{d\mu_{0}}$ be the probability density of parameters $x$ at time $t$ [^13] Formally, the Radon-Nikodym derivative of $\mu_{t}$ w.r.t. $\mu_{0}$. This quantity exists if $\mu_{t}$ is absolutely continuous w.r.t. $\mu_{0}$. We can decompose the posterior probability of behaviorally distinct regions of parameter space, such as regions of low loss with differing degrees of generalization, as follows.
Let $A\subset\mathbb{R}^{N}$ and $B\subset\mathbb{R}^{N}$ be two disjoint regions of parameter space, both with consistently low training loss,141414i.e. contained in a low-loss manifold (Benton et al., 2021) but perhaps distinguished by their performance on a held-out test set. The probability that training will yield an element of $A$ can be decomposed as
[…]
and the log probability ratio is
The strong volume hypothesis would imply that the densities of $A$ and $B$ would be a function only of their training loss, so that the density ratio term would be zero at late times.
A restricted form of the volume hypothesis, for suitable choices of $A$ and $B$, is as follows: Even at the end of training, the volume ratio term in Eq.23 should be larger than the density ratio term.161616 Of course, if the networks in $A$ and the networks in $B$ differ significantly in terms of their performance on the training set, the density ratio term must become very large as $t\rightarrow\infty$, since a well-tuned optimizer is guaranteed to bring the loss close to a local minimum. This is why we require $A$ and $B$ to be low-training-loss.
Appendix C Naive estimator gives a lower bound on local volume
Underestimation problem.
with equality if and only if the log-estimator is constant.
This gap is especially large when the variance of the log-estimator is much larger than 1. For example, if the log-estimator is normally distributed with standard deviation $\sigma$, [^17] As shown theoretically in Appendix A.1 and confirmed in our empirical results, the standard deviation in our case is in fact of order $n$, so this gap can be expected to be significant. the gap is $\sigma^{2}/2$. In these cases, taking more samples will not necessarily fix the problem, because the vast majority of samples will underestimate by a similar amount. A more involved toy model of this phenomenon is given in Appendix A.2.
Another way to phrase this same result is that the estimator for the volume itself (rather than its log) has a very heavy right tail.
Smooth maximum.
In practice, $n$ will be extremely large, ranging from $10^{6}$ to $10^{12}$ parameters. It is therefore worth considering the behavior of our estimator as in the large-$n$ limit. First note that
$\mathbb{E}[\log\widehat{\mathrm{vol}(S)}]\propto\mathbb{E}[\log\sum_{i=1}^{k}\exp\big{(}n\log r({\bm{u}}_{i})\big{)}].$ (25)
$\lim_{n\rightarrow\infty}\frac{1}{n}\log\sum_{i=1}^{k}\exp\big{(}nx_{i}\big{)}=\max(\{x_{1},\ldots,x_{k}\}).$ (26)
This suggests that, in the large-$n$ limit, the normalized log volume estimate $\frac{1}{n}\mathbb{E}[\log\widehat{\mathrm{vol}(S)}]$ will be close to the maximum of our log-radius samples. Empirically, we find that this is already very nearly true for tiny networks of a few thousand parameters (Figure 1).
Markov’s inequality.
That is, the probability that we overestimate the true volume by $m>0$ orders of magnitude is at most one in $10^{m}$. As we expect theoretically and confirm empirically, our Monte Carlo samples for the volume vary over thousands or millions of orders of magnitude. Therefore, overestimating by $m=100$ orders of magnitude would be an imperceptible overestimate, and is extraordinarily unlikely (probability below $10^{-100}$). For this reason, we can treat our naive estimator as a very confident lower bound on the true local volume.
Appendix D Results on ConvNeXt training data
We report here results on ConvNeXt for local KL volume computed over the clean split of the training data (rather than the validation data as in Section 6).
Figure 9: Local volume decrease while training ConvNeXt V2 Atto